OpenStudy (anonymous):

x-y+1=0 find the slope. please explain how you did it so i can learn.

OpenStudy (anonymous):

put it in slope intercept form and get y = mx+b y = x+1 so your m is your slope(basically the number next to the x) in this case its 1 so your slope is 1

OpenStudy (anonymous):

can you explain in a more detailed way? how did you know how to move the equation around like that? and what happened to the rest of the numbers?

OpenStudy (anonymous):

slope intercept form is y=mx+b you have x-y+1=0 so move the y to the right side and get x+1 = y its the same as y = x+1 y=x+1 is in the form y=mx+b so the m is the slope, which is also the number next to x so your slope if the number next to x of y=x+1 so the number next to x is 1 therefore your slope is 1